# Large Language Model Supply Chain: A Research Agenda

### v2

@@ 2.2. Beyond OSS Supply Chain: The New Frontier
Compositional Differences. The OSS and LLM supply chains exhibit fundamental differences across various components, reflecting their distinct natures and requirements. In terms of resources, OSS primarily relies on source code, documentation, repository commits, and package metadata, whereas LLM supply chains are built upon large-scale datasets, pre-trained models, and model-specific metadata like model cards. This divergence extends to development tools, with OSS utilizing traditional software development tools such as version control systems and IDEs, while LLM development necessitates AI-specific frameworks, specialized hardware, and tools for advanced tasks like retrieval-augmented generation (RAG) and model quantization. The core products of these supply chains also differ significantly: OSS typically yields software applications, libraries, and APIs, while LLM supply chains produce pre-trained models, model APIs, and AI agents. Distribution channels follow suit, with OSS primarily leveraging code repositories and package managers, in contrast to LLM products often being distributed through model hubs and AIOps cloud platforms. Deployment and maintenance practices further underscore these differences, as OSS deployment involves software installation and containerization, whereas LLM deployment focuses on model serving and API integration. Similarly, OSS maintenance centers on feature updates and bug fixes, while LLM maintenance encompasses model fine-tuning, retraining, and monitoring for model drift. Lastly, quality assurance methods diverge, with OSS relying on traditional software testing approaches like unit testing and code reviews, in contrast to LLM quality assurance, which involves model-specific evaluations such as bias assessment and AI red teaming.
@@
  • SE Perspective: From the SE standpoint, LLM supply chains inherit issues related to code clones, third-party library dependencies, and open-source license compliance. However, they also introduce new challenges specific to AI models. These include model clone detection, model dependency management, and risk propagation related to hallucination, bias, and toxicity. Novel SE research topics for LLMs also encompass model license compliance, model patch migration, prompt engineering best practices, and model interpretability.
  • S&P Perspective: LLM supply chains inherit several security concerns from OSS, including source code vulnerabilities, dependency and vulnerability propagation, and supply chain poisoning attacks. However, LLMs introduce novel security challenges such as dataset poisoning, model deserialization attacks (e.g., backdoors), prompt injection, and jailbreak attacks. Additionally, LLMs raise concerns about vulnerable autonomous agents and various privacy-related attacks like model inversion and membership inference.
In conclusion, while LLM supply chains share some commonalities with OSS supply chains, they present a unique set of challenges and research opportunities. These span the entire lifecycle of LLMs, from data collection and model training to deployment and governance. The intertwined nature of OSS and LLM supply chains necessitates an interdisciplinary approach to research, combining insights from responsible AI, security, software engineering, ethics, and law to address both inherited and novel challenges in this evolving landscape.
@@ 2.3. Beyond ML/DL Supply Chain: Scaling Complexity
Infrastructure Layer. At the infrastructure layer, LLMs demand unprecedented computational resources, necessitating the extensive use of distributed training frameworks, which introduces new security risks and challenges (Valdez Ladd, 2024) that were less prevalent in traditional ML/DL pipelines. The LLM era has also spawned a plethora of new development tools and features, from vector retrieval systems (Han et al., 2023; facebookresearch, 2024) to inference optimization (ggerganov, 2024) and deployment platforms (mlc-ai, 2024). The security characteristics and software engineering best practices for these rapidly evolving toolchains remain understudied, potentially propagating vulnerabilities throughout the LLM supply chain. Moreover, datasets in the LLM era have gained unprecedented importance, raising critical issues such as privacy concerns (Carlini et al., 2021; Kim et al., 2024), bias mitigation (Ferrara, 2023; Kirk et al., 2021), and license compatibility (Karamolegkou et al., 2023; Katzy et al., 2024a) - topics that were seldom discussed in the ML/DL era.
@@
Foundation Model Layer. In the foundation model layer, the reuse of pre-trained LLMs is far more common and consequential than in traditional ML/DL approaches. This shift introduces a critical challenge: the potential for cascading effects, where issues such as privacy violations (Carlini et al., 2021; Lukas et al., 2023), biases (Gallegos et al., 2023; Kirk et al., 2021), or security vulnerabilities (Yi et al., 2024; Zhou et al., 2024) can propagate from foundation models to a multitude of downstream models and applications. The severity and breadth of this impact far exceed what was typically encountered in ML/DL supply chains, necessitating the need for more rigorous and comprehensive testing and evaluation protocols. Comprehensive testing now includes functionality and performance evaluation (Dong et al., 2023; Sun et al., 2024a), assessing the model’s capabilities across a wide range of tasks and domains; reliability and truthfulness evaluation (Sun et al., 2024c; Li et al., 2023a; Lin et al., 2021), measuring the model’s consistency and tendency to generate factual information; and safety and ethical compliance evaluation (Chao et al., 2024; Gehman et al., 2020; Huang et al., 2023), examining the model’s adherence to ethical guidelines and safety constraints. Furthermore, the proliferation of model hubs for sharing and distributing pre-trained models introduces new security considerations. The increased reliance on shared models in the LLM era has led to growing scrutiny of model hub security. Potential vulnerabilities in these platforms could lead to model or dataset poisoning (Zhao et al., 2024c), compromising model integrity on a scale unseen in the ML/DL era.
@@
Downstream Application Ecosystems. The downstream application layer for LLMs introduces unprecedented complexity. LLMs are being integrated into a wide array of novel application forms, such as chatbots and LLM agents, which are significantly more sophisticated than traditional ML/DL applications. For instance, LLM agents often incorporate components like Retrieval-Augmented Generation (RAG) systems and code execution sandboxes (Li et al., 2024f; He et al., 2024c), expanding the attack surface and introducing security challenges that were non-existent in the ML/DL era (OWASP, 2024b). LLM-powered chatbots face new threat models that are not relevant in traditional ML/DL applications. These include prompt injection attacks (Liu et al., 2023a; Greshake et al., 2023), where malicious users attempt to manipulate the model’s behavior through carefully crafted inputs; function calling attacks (OWASP, 2024b), exploiting the model’s ability to call external functions, potentially leading to unauthorized actions; and prompt leakage (Su et al., 2024; Hou et al., 2024a), the unintended disclosure of sensitive information embedded in prompts or model responses. As more sophisticated LLM agents emerge, additional concerns arise, such as agent autonomy escalation (OWASP, 2024b), where agents potentially exceed their intended level of autonomy, leading to unpredictable or undesired behaviors. These challenges highlight the need for robust safety measures, continuous monitoring, and ethical guidelines in LLM-powered applications, considerations that were less critical in traditional ML/DL deployments.
@@ 2.4. Beyond Responsible/Trustworthy AI: Cascading Impacts
Concept & Research Scope. Responsible/trustworthy AI has traditionally focused on ethical considerations, fairness, transparency, and accountability in AI systems (Kenthapadi et al., 2023; Sun et al., 2024c; Kaur et al., 2022), primarily concerning itself with the societal impact and ethical implications of individual models or applications. In contrast, the LLM supply chain perspective, as explored in this paper, encompasses a broader view that includes both S&P and SE aspects, which can be seen as a superset of responsible/trustworthy AI concepts. While responsible/trustworthy AI typically concentrates on the properties of individual models (Sun et al., 2024c), the LLM supply chain framework examines the cascading effects between models and throughout the entire LLM ecosystem. It considers how vulnerabilities, biases, or other issues can propagate through interconnected components, affecting multiple models and applications downstream. Furthermore, the LLM supply chain perspective incorporates aspects that are often overlooked in responsible/trustworthy AI discussions, such as the impact of infrastructure security on downstream models and applications. This broader view considers the entire technological stack supporting LLMs, from hardware to software infrastructure, and how vulnerabilities at any level can affect the entire ecosystem. For instance, one vulnerability in the infrastructure could potentially compromise multiple models and applications simultaneously (Record, 2024; Nate Nelson, 2024), leading to data leaks, model theft, or unauthorized access. Traditional responsible AI frameworks may not fully address these infrastructure-level risks and their far-reaching consequences. By encompassing both the foundational aspects of responsible AI and the wider considerations of system-level interactions and security, the LLM supply chain provides a more holistic perspective to understanding and mitigating risks in large-scale LLM systems.
@@ 4.1. Infrastructure: Computational Resources
At the heart of model infrastructure are the computational resources, which include hardware such as Graphics Processing Units (GPUs), Tensor Processing Units (TPUs), and other specialized AI accelerators, as well as high-performance CPUs and memory systems. These hardware components are complemented by software frameworks, distributed computing systems, and cloud computing platforms that enable efficient utilization and management of these resources. This vision, encompassing both SE and S&P perspectives, aims to establish a computational infrastructure that is resilient, scalable, and inherently secure.
@@
From an SE perspective, the cornerstone of this vision is the seamless integration of robust SE practices into the management and optimization of computational resources for LLM development. It emphasizes the adoption of Infrastructure as Code (IaC) principles to enable version-controlled, reproducible, and scalable infrastructure setups. Emphasis is placed on creating highly scalable and flexible architectures that can efficiently handle the ever-increasing demands of LLM training and inference. This includes the development of intelligent resource allocation systems and orchestration tools that dynamically adjust to workload variations, ensuring optimal utilization of GPUs, TPUs, and other specialized AI accelerators. The vision also encompasses the implementation of advanced monitoring and diagnostics tools, enabling real-time performance tracking and predictive maintenance to minimize downtime and maximize resource efficiency.
@@
From an S&P perspective, our vision champions a comprehensive “defense-in-depth” approach to securing the computational resources in the LLM supply chain. This holistic strategy addresses vulnerabilities at every level, from hardware to cloud infrastructure and software frameworks. At the hardware level, we envision the development of robust security measures to mitigate risks associated with over-reliance on specific vendors and hardware vulnerabilities. For cloud infrastructure, our vision emphasizes the development of advanced security frameworks specifically tailored for AI workflows. This includes fine-grained access control mechanisms, secure multi-tenant architectures, and end-to-end encryption for data in transit and at rest. In the realm of software frameworks, we envision the integration of security-by-design principles into distributed training systems. This includes rigorous input validation, secure serialization protocols, and sandboxed execution environments for user-defined functions.
@@
Current State. The computational resources landscape in the LLM supply chain is characterized by a complex interplay of hardware advancements, cloud infrastructure, and software frameworks. NVIDIA’s GPUs dominate the market (Dorian Maillard, 2024), leading to potential supply chain vulnerabilities due to over-reliance on a single vendor (Eclypsium, 2024; Ionut Ilascu, 2021; Tyler Sorensen and Heidy Khlaaf, 2024). As an example of hardware vulnerabilities in the LLM supply chain, NVIDIA addressed a critical vulnerability (CVE-2023-31029) in NVIDIA DGX A100 baseboard management controllers (BMCs) (NVIDIA, 2024), which allows unauthenticated attackers to execute arbitrary code and tamper with data remotely, potentially enabling the installation of persistent backdoors in the BMC firmware (Eclypsium, 2024). Such backdoors could control all software executed on the server, including LLM models, training datasets, and model outputs, highlighting the severe implications of hardware vulnerabilities in the LLM supply chain. The trend towards larger and more complex models has led to the widespread adoption of distributed computing systems and specialized AI cloud services. Major cloud providers like AWS, Google Cloud, and Azure have become integral to the LLM infrastructure, offering specialized AI services that, while enabling scalability, introduce new security risks (Jair Ribeiro, 2024; Shaerul Haque J., 2024). On the software front, distributed training frameworks such as Horovod and PyTorch Distributed have enabled efficient large-scale model training but have also introduced new attack vectors. Recently, a critical vulnerability (CVE-2024-5480) has been discovered in PyTorch’s distributed RPC system. This vulnerability, stemming from insufficient input validation, could allow remote code execution when worker nodes serialize and send PythonUDFs (User Defined Functions) to another node without proper restrictions. This flaw potentially enables attackers to execute arbitrary code on affected systems, leading to unauthorized access to sensitive data, manipulation of AI model training processes, and compromise of the entire distributed computing environment (Valdez Ladd, 2024).
@@
Challenge I [S&P Perspective]: Hardware Vulnerabilities and Attacks. Reliance on proprietary hardware solutions introduces vendor lock-in risks and supply chain vulnerabilities. Ensuring compatibility, optimizations, and security across the hardware stack is challenging, particularly when integrating components from diverse suppliers. The proprietary nature of many hardware designs and firmware creates a “black box” scenario, where potential vulnerabilities or backdoors (Zhu et al., 2017; Yang et al., 2016; Gohil et al., 2022a) may remain undetected. Moreover, side-channel attacks pose a significant threat, particularly in the context of GPUs. Due to their parallel processing nature and shared resources, GPUs can leak sensitive information, including model parameters (Naghibijouybari et al., 2018), through timing differences (Jiang et al., 2019), power consumption patterns (Taneja et al., 2023), or electromagnetic emissions (Zhan et al., 2022). The complexity of GPU architectures and their proprietary nature further complicate the implementation of comprehensive security measures against such attacks. These challenges provide some opportunities and future research directions:
@@
  • Opportunity: Secure Hardware Design. The unique requirements of LLM training and inference present an opportunity to rethink hardware design with security as a fundamental principle. This could involve the development of new GPU architectures that incorporate security features at the hardware level, such as dedicated secure enclaves for sensitive computations (Zhu et al., 2020; Mai et al., 2023). Additionally, there’s potential for creating AI-specific hardware accelerators with built-in security measures (Na et al., 2021), designed to resist side-channel attacks and other hardware-level threats. Such secure-by-design hardware could provide a strong foundation for building trustworthy AI, ensuring the integrity and confidentiality of LLM operations from the hardware level up.
@@
  • Opportunity: Hardware Trojan/Backdoor Mitigation. The complexity of hardware-level threats opens up opportunities for developing sophisticated detection and mitigation mechanisms. Current approaches primarily fall into two categories (Gohil et al., 2022a): logic testing and side-channel analysis. Logic testing involves applying input patterns to the chip and monitoring output responses to detect deviations from expected behavior, crucially relying on Trojan-free (or golden) chips as a reference (Chakraborty et al., 2009; Pan and Mishra, 2021; Vashistha et al., 2022). Side-channel analysis, in contrast, focuses on monitoring physical characteristics such as power consumption (Zhao et al., 2020; Salmani and Tehranipoor, 2012), path delay (Jin and Makris, 2008), or electromagnetic emissions (Agrawal et al., 2007; He et al., 2017) to identify anomalies that might indicate the presence of a Trojan. Both methods face challenges such as requiring numerous input patterns (Valdez Ladd, 2024; Pan and Mishra, 2021), low detection rates (Valdez Ladd, 2024; Chakraborty et al., 2009), and, for logic testing, reliance on hard-to-obtain golden chips (He et al., 2017; Vashistha et al., 2022). To address these challenges, future research could focus on developing more efficient and accurate detection methods that reduce dependence on golden chips and minimize the number of required input patterns. Advanced AI techniques, particularly reinforcement learning (Gohil et al., 2022b) and LLM (Chaudhuri et al., 2024; Kokolakis et al., 2024), show potential for improving detection accuracy and efficiency. Future research in these areas may lead to improvements in hardware security practices, contributing to the integrity and trustworthiness of LLM systems from the hardware level up.
@@
Challenge II [S&P Perspective]: Distributed Training Risks. The advent of LLMs has necessitated the adoption of distributed training approaches, introducing a complex array of security challenges and supply chain risks. This distributed infrastructure, while enabling unprecedented model capabilities, significantly expands the attack surface, with each node representing a potential vulnerability (Lyu et al., 2020). Node failures, whether due to hardware malfunctions, network issues, or targeted attacks, pose critical risks to training continuity and data integrity (Liu et al., 2023d; Verbraeken et al., 2020). Moreover, the complexity of distributed training frameworks, such as PyTorch Distributed, Horovod, and TensorFlow Distributed, also expand the attack surface significantly. First, there is the potential for inadequate encryption or authentication mechanisms, which could expose critical components like gradient updates and model parameters to interception or tampering by malicious actors. Second, the complexity of these systems can sometimes lead to misconfigurations, such as exposed APIs or insufficient access controls (Valdez Ladd, 2024; Record, 2024), which may create additional security vulnerabilities.
@@
Challenge III [SE Perspective]: Computational Resource Management and Optimization. The computational resources in the Large Language Model (LLM) supply chain present unique challenges and opportunities from an SE perspective. One significant challenge is the need for efficient resource allocation and optimization in distributed training environments. As LLMs grow in size and complexity, traditional SE practices struggle to manage the massive parallelism and data flow required for training. This necessitates the development of novel scheduling algorithms and load-balancing techniques that can dynamically adapt to the varying computational demands of different training phases. Another challenge lies in ensuring the reproducibility and version control of training environments, as subtle differences in hardware configurations or software dependencies can lead to inconsistent model behaviors.
  • Opportunity: Advancing IaC and Intelligent Compute Scheduling. The challenges in LLM resource management present an opportunity to advance IaC practices and develop intelligent compute scheduling systems tailored for AI workflows. This involves creating sophisticated IaC tools that can dynamically provision and manage the complex, heterogeneous infrastructure required for LLM training. Such tools would enable version-controlled, reproducible environment setups across different cloud providers and on-premises systems, ensuring consistency in training results. Furthermore, there is an opportunity to develop AI-driven scheduling algorithms that optimize resource allocation in real-time. These schedulers could leverage machine learning techniques to predict resource demands, automatically adjust to workload variations, and efficiently distribute tasks across available GPUs, TPUs, and other AI accelerators.
  • Opportunity: Advancing Green SE Practices for Energy-efficient Training. The increasing focus on the environmental impact of AI training presents an opportunity to implement green SE practices specifically for LLM development. This involves creating methodologies and tools to optimize energy consumption without compromising model performance. Potential areas of innovation include the development of energy-aware scheduling algorithms that prioritize workloads based on both performance requirements and energy efficiency. Software engineers could focus on creating more efficient model architectures and training algorithms that reduce computational requirements while maintaining model quality.

### v3

@@ 2.2. Beyond OSS Supply Chain: The New Frontier
Compositional Differences. The compositional differences between the LLM supply chain and the OSS supply chain contribute to the divergence in their supply chains, resulting in the LLM supply chain inheriting some research topics from the OSS supply chain while also introducing novel challenges. In terms of resources, OSS primarily relies on source code, documentation, repository commits, and package metadata, whereas LLM supply chains are built upon large-scale datasets, pre-trained models, prompts, and model-specific metadata like model cards. Besides, the OSS utilizes traditional software development tools, while LLM development necessitates AI-specific frameworks, specialized hardware, and tools for advanced tasks like retrieval-augmented generation (RAG), model quantization, and model merging. The distribution channels of these supply chains also differ significantly: OSS primarily leverages code repositories and package managers, in contrast to LLM products often being distributed through model hubs and AIOps platforms. Similarly, OSS maintenance centers on feature updates and bug fixes, while LLM maintenance encompasses model fine-tuning, retraining, and monitoring for model drift. Lastly, quality assurance methods diverge, with OSS relying on traditional software testing approaches like unit testing and code reviews, in contrast to LLM quality assurance, which involves model-specific evaluations such as bias assessment and AI red teaming.
@@
  • SE Perspective: From the SE standpoint, LLM supply chains inherit issues related to code clones, third-party library dependencies, and open-source license compliance. However, they also introduce new challenges specific to pre-trained models, which include model clone detection, model dependency management and model license compliance.
  • S&P Perspective: LLM supply chains inherit several security concerns from OSS, including source code vulnerabilities, dependency and vulnerability propagation, and supply chain poisoning attacks. However, LLMs introduce novel security challenges such as dataset poisoning, model deserialization attacks (e.g., backdoors), and code hallucination. Additionally, LLMs raise concerns about prompt injection, jailbreak attacks, and vulnerable autonomous agents.
@@ 2.3. Beyond ML/DL Supply Chain: Scaling Complexity
Infrastructure Layer. At the infrastructure layer, LLMs demand unprecedented computational resources, necessitating the extensive use of distributed training frameworks, which introduces new security risks and challenges (Valdez Ladd, 2024) that were less prevalent in traditional ML/DL pipelines. The LLM era has also spawned a plethora of new development tools and features, from vector retrieval systems (Han et al., 2023; facebookresearch, 2024) to inference optimization (ggerganov, 2024) and deployment platforms (mlc-ai, 2024). The security characteristics and software engineering best practices for these rapidly evolving toolchains remain understudied, potentially propagating vulnerabilities throughout the LLM supply chain. Moreover, datasets in the LLM era have gained unprecedented importance, raising critical issues such as privacy concerns (Carlini et al., 2021; Kim et al., 2023), bias mitigation (Ferrara, 2023; Kirk et al., 2021), and license compatibility (Karamolegkou et al., 2023; Katzy et al., 2024a) - topics that were seldom discussed in the ML/DL era.
@@
Foundation Model Layer. In the foundation model layer, the reuse of pre-trained LLMs is far more common and consequential than in traditional ML/DL approaches. This shift introduces a critical challenge: the potential for cascading effects, where issues such as privacy violations (Carlini et al., 2021; Lukas et al., 2023), biases (Gallegos et al., 2023; Kirk et al., 2021), or security vulnerabilities (Yi et al., 2024; Zhou et al., 2024) can propagate from foundation models to a multitude of downstream models and applications. The severity and breadth of this impact far exceed what was typically encountered in ML/DL supply chains, necessitating the need for more rigorous and comprehensive testing and evaluation protocols. Comprehensive testing now includes functionality and performance evaluation (Dong et al., 2023; Sun et al., 2024a), assessing the model’s capabilities across a wide range of tasks and domains; reliability and truthfulness evaluation (Sun et al., 2024c; Li et al., 2023a; Lin et al., 2022), measuring the model’s consistency and tendency to generate factual information; and safety and ethical compliance evaluation (Chao et al., 2024; Gehman et al., 2020; Huang et al., 2023), examining the model’s adherence to ethical guidelines and safety constraints. Furthermore, the proliferation of model hubs for sharing and distributing pre-trained models introduces new security considerations. The increased reliance on shared models in the LLM era has led to growing scrutiny of model hub security. Potential vulnerabilities in these platforms could lead to model or dataset poisoning (Zhao et al., 2024b), compromising model integrity on a scale unseen in the ML/DL era.
@@
Downstream Application Ecosystems. The downstream application layer for LLMs introduces unprecedented complexity. LLMs are being integrated into a wide array of novel application forms, such as chatbots and LLM agents, which are significantly more sophisticated than traditional ML/DL applications. For instance, LLM agents often incorporate components like retrieval-augmented generation (RAG) systems and code execution sandboxes (Li et al., 2024e; He et al., 2024c), expanding the attack surface and introducing security challenges that were non-existent in the ML/DL era (OWASP, 2024b). LLM-powered chatbots face new threat models that are not relevant in traditional ML/DL applications. These include prompt injection attacks (Liu et al., 2023a; Abdelnabi et al., 2023), where malicious users attempt to manipulate the model’s behavior through carefully crafted inputs; function calling attacks (OWASP, 2024b), exploiting the model’s ability to call external functions, potentially leading to unauthorized actions; and prompt leakage (Su et al., 2024; Hou et al., 2024a), the unintended disclosure of sensitive information embedded in prompts or model responses. As more sophisticated LLM agents emerge, additional concerns arise, such as agent autonomy escalation (OWASP, 2024b), where agents potentially exceed their intended level of autonomy, leading to unpredictable or undesired behaviors. These challenges highlight the need for robust safety measures, continuous monitoring, and ethical guidelines in LLM-powered applications, considerations that were less critical in traditional ML/DL deployments.
@@ 2.4. Beyond Responsible/Trustworthy AI: Cascading Impacts
Concept & Research Scope. Responsible/trustworthy AI has traditionally focused on ethical considerations, fairness, transparency, and accountability in AI systems (Kenthapadi et al., 2023; Sun et al., 2024c; Kaur et al., 2023), primarily concerning itself with the societal impact and ethical implications of individual models or applications. In contrast, the LLM supply chain perspective, as explored in this paper, encompasses a broader view that includes both S&P and SE aspects, which can be seen as a superset of responsible/trustworthy AI concepts. While responsible/trustworthy AI typically concentrates on the properties of individual models (Sun et al., 2024c), the LLM supply chain framework examines the cascading effects between models and throughout the entire LLM ecosystem. It considers how vulnerabilities, biases, or other issues can propagate through interconnected components, affecting multiple models and applications downstream. Furthermore, the LLM supply chain perspective incorporates aspects that are often overlooked in responsible/trustworthy AI discussions, such as the impact of infrastructure security on downstream models and applications. This broader view considers the entire technological stack supporting LLMs, from hardware to software infrastructure, and how vulnerabilities at any level can affect the entire ecosystem. For instance, one vulnerability in the infrastructure could potentially compromise multiple models and applications simultaneously (Record, 2024; Nate Nelson, 2024), leading to data leaks, model theft, or unauthorized access. Traditional responsible AI frameworks may not fully address these infrastructure-level risks and their far-reaching consequences. By encompassing both the foundational aspects of responsible AI and the wider considerations of system-level interactions and security, the LLM supply chain provides a more holistic perspective to understanding and mitigating risks in large-scale LLM systems.
@@ 4.1. Infrastructure: Computational Resources
At the heart of model infrastructure are the computational resources, which include hardware such as graphics processing units (GPUs), tensor processing units (TPUs), and other specialized AI accelerators. These hardware components are complemented by software frameworks, distributed training frameworks, and cloud computing platforms that enable efficient utilization and management of these resources. Our vision, encompassing both SE and S&P perspectives, aims to establish a computational infrastructure that is resilient, scalable, and inherently secure.
@@
From an SE perspective, our vision is the seamless integration of robust SE practices into the management and optimization of computational resources for LLM development. It emphasizes the adoption of Infrastructure as Code (IaC) principles to enable version-controlled, reproducible, and scalable infrastructure setups. Emphasis is placed on creating highly scalable and flexible architectures that can efficiently handle the ever-increasing demands of LLM training and inference. This includes the development of intelligent resource allocation systems and orchestration tools that dynamically adjust to workload variations, ensuring optimal utilization of GPUs, TPUs, and other specialized AI accelerators. The vision also encompasses the implementation of advanced monitoring and diagnostics tools, enabling real-time performance tracking and predictive maintenance to minimize downtime and maximize resource efficiency.
@@
From an S&P perspective, our vision champions a comprehensive “defense-in-depth” approach to securing the computational resources in the LLM supply chain. This holistic strategy addresses vulnerabilities at every level, from hardware to cloud infrastructure and software frameworks. At the hardware level, we envision the development of robust security measures to mitigate risks associated with over-reliance on specific vendors and hardware vulnerabilities. For cloud infrastructure, our vision emphasizes the development of advanced security frameworks specifically tailored for AI workflows. This includes fine-grained access control mechanisms and secure multi-tenant architectures. In the realm of software frameworks, we envision the integration of security-by-design principles into distributed training systems. This includes rigorous input validation, secure serialization protocols, and sandboxed execution environments for user-defined functions.
@@
Current State. The computational resources landscape in the LLM supply chain is characterized by a complex interplay of hardware advancements, cloud infrastructure, and software frameworks. NVIDIA’s GPUs dominate the market (Dorian Maillard, 2024), leading to potential supply chain vulnerabilities due to over-reliance on a single vendor (Eclypsium, 2024; Ionut Ilascu, 2021; Sorensen and Khlaaf, 2024). As an example of hardware vulnerabilities in the LLM supply chain, NVIDIA addressed a critical vulnerability (CVE-2023-31029) in NVIDIA DGX A100 baseboard management controllers (BMCs) (NVIDIA, 2024), which allows unauthenticated attackers to execute arbitrary code and tamper with data remotely, potentially enabling the installation of persistent backdoors in the BMC firmware (Eclypsium, 2024). Such backdoors could control all software executed on the server, including LLM models, training datasets, and model outputs, highlighting the severe implications of hardware vulnerabilities in the LLM supply chain. The trend towards larger and more complex models has led to the widespread adoption of distributed computing systems and specialized AI cloud services. Major cloud providers like AWS, Google Cloud, and Azure have become integral to the LLM infrastructure, offering specialized AI services that, while enabling scalability, introduce new security risks (Jair Ribeiro, 2024; Shaerul Haque J., 2024). On the software front, distributed training frameworks such as Horovod and PyTorch Distributed have enabled efficient large-scale model training but have also introduced new attack vectors. Recently, a critical vulnerability (CVE-2024-5480) has been discovered in PyTorch’s distributed RPC system. This vulnerability, stemming from insufficient input validation, could allow remote code execution when worker nodes serialize and send PythonUDFs (User Defined Functions) to another node without proper restrictions. This flaw potentially enables attackers to execute arbitrary code on affected systems, leading to unauthorized access to sensitive data, manipulation of model training processes, and compromise of the entire distributed computing environment (Valdez Ladd, 2024).
@@
Challenge I [S&P Perspective]: Hardware Vulnerabilities and Attacks. Reliance on proprietary hardware solutions introduces vendor lock-in risks and supply chain vulnerabilities. Ensuring compatibility, optimizations, and security across the hardware stack is challenging, particularly when integrating components from diverse suppliers. The proprietary nature of many hardware designs and firmware creates a “black box” scenario, where potential vulnerabilities or backdoors (Zhu et al., 2017; Yang et al., 2016; Gohil et al., 2022) may remain undetected. Moreover, side-channel attacks pose a significant threat, particularly in the context of GPUs. Due to their parallel processing nature and shared resources, GPUs can leak sensitive information, including model parameters (Naghibijouybari et al., 2018), through timing differences (Jiang et al., 2020), power consumption patterns (Taneja et al., 2023), or electromagnetic emissions (Zhan et al., 2022). The complexity of GPU architectures and their proprietary nature further complicate the implementation of comprehensive security measures against such attacks. These challenges provide some opportunities and future research directions:
@@
  • Opportunity: Secure Hardware Design. The unique requirements of LLM training and inference present an opportunity to rethink hardware design with security as a fundamental principle. This could involve incorporating security features at the hardware level (Zhu et al., 2020; Mai et al., 2023), designed to resist side-channel attacks and other hardware-level threats. Such secure-by-design hardware could provide a strong foundation for building trustworthy AI, ensuring the integrity and confidentiality of LLM operations from the hardware level up.
@@
  • Opportunity: Hardware Trojan/Backdoor Mitigation. The complexity of hardware-level threats opens up opportunities for developing sophisticated detection and mitigation mechanisms. Current approaches primarily fall into two categories (Gohil et al., 2022): logic testing and side-channel analysis. Logic testing involves applying input patterns to the chip and monitoring output responses to detect deviations from expected behavior, crucially relying on Trojan-free (or golden) chips as a reference (Chakraborty et al., 2009; Pan and Mishra, 2021; Vashistha et al., 2022). Side-channel analysis, in contrast, focuses on monitoring physical characteristics such as power consumption (Zhao et al., 2020; Salmani and Tehranipoor, 2012), path delay (Jin and Makris, 2008), or electromagnetic emissions (Agrawal et al., 2007; He et al., 2017) to identify anomalies that might indicate the presence of a Trojan. Both methods face challenges such as requiring numerous input patterns (Valdez Ladd, 2024; Pan and Mishra, 2021), low detection rates (Valdez Ladd, 2024; Chakraborty et al., 2009), and, for logic testing, reliance on hard-to-obtain golden chips (He et al., 2017; Vashistha et al., 2022). To address these challenges, future research could focus on developing more efficient and accurate detection methods that reduce dependence on golden chips and minimize the number of required input patterns. Advanced AI techniques, particularly reinforcement learning (Gohil et al., 2024) and LLM (Chaudhuri et al., 2024; Kokolakis et al., 2024), show potential for improving detection accuracy and efficiency. Future research in these areas may lead to improvements in hardware security practices, contributing to the integrity and trustworthiness of LLM systems from the hardware level up.
@@
Challenge II [S&P Perspective]: Distributed Training Risks. The advent of LLMs has necessitated the adoption of distributed training approaches, introducing a complex array of security challenges and supply chain risks. This distributed infrastructure, while enabling unprecedented model capabilities, significantly expands the attack surface, with each node representing a potential vulnerability (Lyu et al., 2020). Node failures, whether due to hardware malfunctions, network issues, or targeted attacks, pose critical risks to training continuity and data integrity (Liu et al., 2023d; Verbraeken et al., 2021). Moreover, the complexity of distributed training frameworks, such as PyTorch Distributed, Horovod, and TensorFlow Distributed, also expand the attack surface significantly. First, there is the potential for inadequate encryption or authentication mechanisms, which could expose critical components like gradient updates and model parameters to interception or tampering by malicious actors. Second, the complexity of these systems can sometimes lead to misconfigurations, such as exposed APIs or insufficient access controls (Valdez Ladd, 2024; Record, 2024), which may create additional security vulnerabilities.
@@
Challenge III [SE Perspective]: Computational Resource Management and Optimization. The computational resources in the LLM supply chain present unique challenges and opportunities from an SE perspective. One significant challenge is the need for efficient resource allocation and optimization in distributed training environments. As LLMs grow in size and complexity, traditional SE practices struggle to manage the massive parallelism and data flow required for training. This necessitates the development of novel scheduling algorithms and load-balancing techniques that can dynamically adapt to the varying computational demands of different training phases. Another challenge lies in ensuring the reproducibility and version control of training environments, as subtle differences in hardware configurations or software dependencies can lead to inconsistent model behaviors.
  • Opportunity: Advancing IaC and Intelligent Compute Scheduling. The challenges in LLM resource management present an opportunity to advance IaC practices and develop intelligent compute scheduling systems tailored for AI workflows (Diaz-De-Arcaya et al., 2024). This involves creating sophisticated IaC tools that can dynamically provision and manage the complex, heterogeneous infrastructure required for LLM training. Such tools would enable version-controlled, reproducible environment setups across different cloud providers and on-premises systems, ensuring consistency in training results.
  • Opportunity: Green SE Practices for LLM Training. The increasing focus on the environmental impact of AI training presents an opportunity to implement green SE practices specifically for LLM development (Shi et al., 2024b). This involves creating methodologies and tools to optimize energy consumption without compromising model performance. Potential areas of innovation include the development of energy-aware scheduling algorithms that prioritize workloads based on both performance requirements and energy efficiency. Software engineers could focus on creating more efficient model architectures and training algorithms that reduce computational requirements while maintaining model quality.